OpenStudy (anonymous):

Find the linearization L(x) of the following function: f(x) = x^3 + x^2 - 4; a=2

OpenStudy (anonymous):

use the following formula \[\Large L(x)=f(a)+f'(a)(x-a)\] does this help ??

OpenStudy (anonymous):

@kathyQ waiting for your reply ..

OpenStudy (anonymous):

ok f(a) =f(2)=(2)^3+(2)^2-4 f(2)=8+4-4=8 f'(x)=3x^2+2x and f'(a) becomes f'(a)=f'(2)=3(2)^2+2(2)=16 just put these values in the above given formula

OpenStudy (anonymous):

sorry, i left to buy some stuff. let me check

OpenStudy (anonymous):

so, the answer would be L(x) = 8 + 16x-16a, right?

OpenStudy (anonymous):

i corrected my answer and plugged in 2 as a